OpenStudy (anonymous):

How do you find the integral of= ʃ(sinx)^2 (cosx)^4 dx

OpenStudy (anonymous):

How do you find the integral of \[\int\limits_{}^{} (sinx)^{2} (cosx)^{4} dx\] knowing that: \[(cosx)^{2} = \frac{ 1 }{ 2 } ( 1 + \cos2x) \]

OpenStudy (anonymous):

and that \[(sinx)^{2} = \frac{ 1 }{ 2 } (1 - \cos2x)\]

OpenStudy (anonymous):

and that \[(\cos2x)^{3} =\cos2x (1 - (\sin2x)^{2})\]

OpenStudy (turingtest):

well if you want to make use of that last identity I think that you should start by changing sin^2x to 1-cos^2x

OpenStudy (turingtest):

it will still be a dreadful mess though

zepdrix (zepdrix):

`Even Powers` of cosine and sine are such a pain. I much prefer to use the Cosine Reduction Formula:\[\Large\bf\sf \int\limits \cos^nx\;dx =\frac{1}{n}\cos^{n-1}x \sin x+\frac{n-1}{n}\int\limits \cos^{n-2}x\;dx\] It reduces the power on cosine by 2 each time ^ But if you prefer Using the Half-Angle Identity over and over to get rid of those squares, that will work also.

OpenStudy (anonymous):

Can you show me the fastest one in a detailed way so i can understand?

OpenStudy (turingtest):

oh I think I have an easier way... let me make sure it pans out

OpenStudy (anonymous):

okk

OpenStudy (turingtest):

nope still a mess :P

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

What do I do then? ç.ç

zepdrix (zepdrix):

\[\Large\bf\int\limits \sin^2x \cos^4x dx\quad=\quad \int\limits \cos^4x\;dx - \int\limits \cos^6x\;dx\] Let,\[\Large\bf I_n\quad=\quad \int\limits \cos^nx\;dx\] Then let's deal with the cos^6 term first,\[\Large\bf I_6\quad=\quad \frac{1}{6}\cos^5x \sin x+\frac{5}{6}I_4\]So that brings us here with the equation we started with:\[\Large\bf \int\limits \cos^4x\;dx -\int\limits \cos^6x\;dx\]\[\Large\bf =\quad I_4-\left(\frac{1}{6}\cos^5x \sin x+\frac{5}{6}I_4\right)\]Now both integrals are in terms of cosine^4 so we can combine them.

zepdrix (zepdrix):

If this is too confusing, especially with my short hand notation D: You can let me know.

zepdrix (zepdrix):

Another thing to keep in mind is.. if we're going to use this reduction formula, your teacher would probably want you to justify where it comes from :d Which also takes a few steps, integration by parts.

OpenStudy (anonymous):

What happened in the first passage? How did you transform it in cos^4(x) - cos^6(x)?

OpenStudy (turingtest):

I'm sure @zepdrix 's way will work. Given the identities you posted, I think they want you to go\[\int\sin^2x\cos^4xdx=\int(1-\cos^2x)\cos^4xdx=\int\cos^4x-\cos^6xdx\\=\int[1+\cos(2x)]^2-[1+\cos(2x)]^3\]expand and yawn...

OpenStudy (anonymous):

oh I got it xD

zepdrix (zepdrix):

Ya if they `gave you` those identities to use, maybe you should go with it that way :)

OpenStudy (anonymous):

@TuringTest the last passage doesn't make it zero?

OpenStudy (anonymous):

Sorry for the ignorance but it's the first time I'm doing this...

OpenStudy (turingtest):

let me make it a bit bigger

OpenStudy (turingtest):

\[\int\sin^2x\cos^4xdx=\int(1-\cos^2x)\cos^4xdx=\int\cos^4x-\cos^6xdx\\=\large\int[1+\cos(2x)]^2-[1+\cos(2x)]^3dx\]

OpenStudy (turingtest):

now that you can read the exponents clearly, do you still think we should get zero?

OpenStudy (anonymous):

no xD

OpenStudy (anonymous):

Thank you for being more detailed ^^

OpenStudy (turingtest):

:)

OpenStudy (anonymous):

\[1 + (\cos2x)^{2} + 2 (\cos2x) - 1 - (\cos2x)^{3} - 3\cos2x - 3 (\cos2x)^{2} \]

OpenStudy (anonymous):

\[\int\limits_{}^{} -2 (\cos2x)^{2} - 2\cos2x - (\cos2x)^{3} dx\]

OpenStudy (turingtest):

I think it's a coefficient of 1 on the cos(2x) term in the middle there

OpenStudy (turingtest):

-1*

OpenStudy (turingtest):

and I totally forgot to write the factors of 1/2 in the original setup, oops

OpenStudy (turingtest):

\[\int\sin^2x\cos^4xdx=\int(1-\cos^2x)\cos^4xdx=\int\cos^4x-\cos^6xdx\\=\large\int\frac14[1+\cos(2x)]^2-\frac18[1+\cos(2x)]^3dx\]

OpenStudy (turingtest):

you were on the right track though... lots of tedious algebra to make mistakes in; very little interesting calculus in this problem

OpenStudy (anonymous):

ah ok. thanks

OpenStudy (anonymous):

Um... should I do it again and show you the result?

OpenStudy (turingtest):

if I were you I'f get the fractions out of the way because they are annoying, and factor out 1/8, giving\[\large\frac18\int2[1+\cos(2x)]^2-[1+\cos(2x)]^3dx\]but that just makes it a little less ugly

OpenStudy (turingtest):

however you prefer. If you want to work the whole thing out on paper, or post each step, it's up to you

OpenStudy (anonymous):

I'll post each step. Thank you for your patience xD

OpenStudy (turingtest):

welcome :)

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 } \int\limits_{}^{} 2( 1 + 2\cos2x + (\cos2x)^{2}) - (1+(\cos2x)^{3}+3\cos2x+3(\cos2x)^{2})\]

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 } \int\limits_{}^{} 2+4\cos2x+2(\cos2x)^{2}-1-(\cos2x)^{3}-3\cos2x-3(\cos2x)^{2}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }\int\limits_{}^{} 1-(\cos2x)^{2}-(\cos2x)^{3}+\cos2x\]

OpenStudy (turingtest):

I agree so far

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }\int\limits_{}^{}1-\cos2x(1-(\sin2x)^{2})-(\cos2x)^{2} + \frac{ 1 }{ 2 }\sin2x\]

OpenStudy (anonymous):

How do i trasform \[(\cos2x)^{2}\] knowing the other formula?

OpenStudy (turingtest):

first of all, where did \(\frac12\sin2x\) come from?

OpenStudy (anonymous):

cos2x

OpenStudy (turingtest):

why 1/2 then?

OpenStudy (anonymous):

I don't know, it's a formula I just took from my book

OpenStudy (turingtest):

books are nice for reference, but when they don't agree with your brain you should decide which one is right let's go back to where we agreed:\[\frac{ 1 }{ 8 }\int 1-(\cos2x)^{2}-(\cos2x)^{3}+\cos2xdx\]

OpenStudy (turingtest):

now we use the two formulas you were given at the beginning of the problem to sub for cos^2(2x) and cos^3(2x)

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }\int\limits_{}^{}1-(\cos2x)^{2}-\cos2x(1-(\sin2x)^{2})+\cos2x\]

OpenStudy (anonymous):

I don't know why I sent the same message D:

OpenStudy (turingtest):

I didn't see it o.0 but this is good so farm but you better deal with that \(\cos^22x\) part repeat the formula you were given to simplify cos^2

OpenStudy (turingtest):

good so far*

OpenStudy (anonymous):

But if I know that \[(cosx)^{2}=\frac{ 1 }{ 2 }(1+\cos2x)\]

OpenStudy (anonymous):

how do I apply it to cos2x?

OpenStudy (turingtest):

the exact same way. just let u=2x

OpenStudy (anonymous):

I mean (cos2x)^2

OpenStudy (turingtest):

are you stuck or just thinking?

OpenStudy (anonymous):

stuck xD sorryyy

OpenStudy (turingtest):

no worries, if you had (cos(x))^2 you would know what to do, right?

OpenStudy (anonymous):

You said that u=2x so should i put it like this? \[\cos(\frac{ u }{ 2 })^2= \frac{ 1 }{ 2 } (1+ cosu)\]

OpenStudy (turingtest):

well you have the right idea, but you mixed up the two ideas we are trying to get \[\cos^2(2x)\]let \(u=2x\) we then have\[\cos^2u=\frac12(1+\cos2u)\]now we just undo our substitution to get...?

OpenStudy (anonymous):

cos\[\cos^{2}2x=\frac{ 1 }{ 2 }(1+\cos4x)\]

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }\int\limits_{}^{}1-\frac{ 1 }{ 2 }(1+\cos4x)-\cos2x(1-(\sin2x )^{2}) +\cos2x\]

OpenStudy (turingtest):

yep, getting pretty ugly, right? now distribute and simplify...

OpenStudy (anonymous):

I'm sleepy xD It's 1:20 in my country \[\frac{ 1 }{ 8 }\int\limits_{}^{}1-\frac{ 1 }{ 2 }-\frac{ 1 }{ 2 }\cos4x-\cos2x+(\cos2x)(\sin2x)^{2}+\cos2x\]

OpenStudy (turingtest):

Well you're doing really well, then. It's only 6:30 here, but I've been up since 5am, so I'm about to call it a day after this long, painful integral

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }\int\limits_{}^{}\frac{ 1 }{ 2}-\frac{ 1 }{ 2 }\cos4x+(\cos2x) (\sin2x)^{2}\]

OpenStudy (anonymous):

yea you're right

OpenStudy (anonymous):

I understood the previous partss so it's ok. Thank again for helping me :D

OpenStudy (turingtest):

I'm willing to finish this if you are, you're almost done here

OpenStudy (anonymous):

*you I'm sleeping xD

OpenStudy (turingtest):

fair enough, the rest can be done with a simple u-substitution. I'm sure that will be apparent after a good night's rest. See ya around!

OpenStudy (anonymous):

Ok! byee!! C: